OpenStudy (anonymous):

Solve: 2(x - 3) + 2(x - 7) = 40

sam (.sam.):

Divide 2 x - 3 + x - 7 = 20 2x-10=20

sam (.sam.):

Then sovle

OpenStudy (anonymous):

???

sam (.sam.):

2x-10=20 2x=30 x=15